e and g differential equations general solutions and uniqe for e
(e) y = x + y; y(0) = 1 (f) y +9y +14y = 0 (g) y= (8) 9 3xy 0222
Answer #1

Bytbly - (917) cb Integrating both side. 4.-3.9% = 49% + fx. te, 77(0321 » Yo - S/S =ė 3)(3-4 >> 320 - 3495 = % + 7x = An36. W . -(6-36-1) -V+3611 Let youa. aufz Bu 3(26-3)+ = 39- o om (202-1)dve alex -62+36 +1 or (1-102) die du 0²36- 12 or, (-(